OpenStudy (anonymous):

i hate factoring can someone lend a hand?? 2y^2-32=0

OpenStudy (anonymous):

y=4.-4will be better....

OpenStudy (saifoo.khan):

u just have to factor right?

OpenStudy (anonymous):

yes but i dont understand what you did dont i need (x plus/minus blank) times (x plus/minus blank)??

OpenStudy (saifoo.khan):

everytime they arent that way.

OpenStudy (anonymous):

so what was the answer?

OpenStudy (saifoo.khan):

2(y^2-16) = 0 2(y-4)(y+4)=0 <---answer.

OpenStudy (anonymous):

can i put the 2 back in if i need to find the solution?

OpenStudy (anonymous):

2y^2-32=0 2(y^2-16)=0 (y^2-4^2)=0 (y-4)(y+4)=0

OpenStudy (saifoo.khan):

yes u can, but its better to write.

OpenStudy (anonymous):

k cool thanks for the help